Circular No. 7829

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS D. K. Lynch, R. J. Rudy, S. Mazuk, and C. Venturini, The Aerospace

Corporation, report 0.8–2.5-μm spectroscopy of V838 Mon on Jan. 28.16 UT: “Using the Aerospace Near-Infrared Imaging Spectrometer on the Lick 3-m telescope, we observed a continuum peaking near 1 μm (photospheric temperature of roughly 3000 K), with significant broad structure between 0.8 and 1.2 μm, and a slightly elevated continuum near 1.6 μm, suggestive of the H^{−} opacity minimum. Emission lines present included H I Paschen and weak Brackett lines, many with strong P-Cyg profiles, indicating an expansion velocity of around 500 km/s. Also present were the Ca II infrared triplet with strong absorption and weak emission, deep CO first-overtone band absorption near 2.3 μm, and a number of the same lines reported by Geballe et al. (IAUC 7796), most of them with P-Cyg profiles; V838 Mon appears to have brightened in the infrared by ∼ 30 percent since Jan. 12.33, and the emission lines have also strengthened.”

N. D. Morrison, University of Toledo, reports a change to higher excitation in the spectrum of V838 Mon since Jan. 26 (IAUC 7812): “CCD échelle spectra (resolving power 26 000; nine 7-nm regions in the yellow and red), were taken with the Ritter Observatory 1-m telescope on Feb. 5.2, 6.1, 8.1, and 9.1 UT by K. S. Bjorkman, A. Miroshnichenko, and myself (and reduced by J. P. Wisniewski). The only stellar features detected in this range were Fe II (531.6 nm), Na I (588.9 and 589.5 nm), Si II (634.7 and 637.1 nm), and Hα, all with P-Cyg profiles. Hα is mainly in emission, has a mean net equivalent width of −3.2 nm, and has electron scattering wings extending at least 3 nm on either side of the line center. Its absorption component has a mean heliocentric blue-edge velocity of −300 km/s, and the mean absorption centroid velocity is −200 km/s. Significant strengthening of the Na I absorption and emission components and weakening of the Si II absorption component took place between Feb. 6 and 8.”

SUPERNOVAE 2002aq AND 2002at A. V. Filippenko, University of California, Berkeley; and A. J. Barth,

California Institute of Technology, report: “Inspection of a CCD spectrum (range 320–1000 nm), obtained on Feb. 14 UT with the Keck-I telescope, reveals that SN 2002aq (IAUC 7813) is of type II, probably within one month past maximum brightness; the Hα line has a well-developed P-Cyg profile. A spectrum of SN 2002at (IAUC 7821) is essentially featureless and quite blue; perhaps the object is a very young type-II supernova.”

2002 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green